OpenStudy (anonymous):

how many pairs of natural numbers are there if the difference of their squares is 60

OpenStudy (shubhamsrg):

a^2 - b^2 = 60 (a+b)(a-b) = (2^2) * 3 * 5 Guess you'll have to take cases here. And compare LHS and RHS ofcorse.

OpenStudy (shubhamsrg):

Note that 60 has 12 factors!

OpenStudy (shubhamsrg):

So you'll have to take 6 cases. 60 = 1*60 = 2*30 = 3*20 = 4*15 = 5*12 = 6*10 Does it help ?

OpenStudy (anonymous):

no i want clearly because the answer is given 2

OpenStudy (sirm3d):

the only possible cases are \((a-b)=2,(a+b)=30\text{ and } (a-b)=6,(a+b)=10\)

OpenStudy (shubhamsrg):

See, you need to compare First thing to keep in mind is a+b will be greater than a-b given that both are natural numbers If a+b is odd, then a-b will also be odd If a+b is even, then a-b will also be even Keep those in mind and compare

OpenStudy (shubhamsrg):

as sire @sirm3d says.

OpenStudy (anonymous):

if 3m = 3x and 3n = 27x then m equals to what?

OpenStudy (shubhamsrg):

m = log(e^x)

OpenStudy (shubhamsrg):

where base of log is e